Dear Richard dear all,

please find below a short summary of the activities done Sunday & today on the mountain related to the AO system. Italian people working at LBTO were:


Arcetri: A. Puglisi, P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni
LBTO: G. Brusa

Total # of person: 8

W unit

WFS optics cleaning:
1) cleanings of the CCD39 camera entrance window removed the detected dust grains.
2) cleaning of the pupil rotator prism. One of the two surfaces of the prism has been cleaned carefully. However about three dimmed spots seen on the pupil images (about 1x1, 2x2 pixel in the CCD) still remain there. Cleaning will be continued tomorrow.
The last devices to be cleaned are: the fast tip tilt mirror and the filters of FW#2

AO control SW

No activity


LBT672a

1) Production of tools for dust protection cover of the central membrane.
2) Test of the elevation slew. the test has been repeated about 20 times. Two aims: one is to check stability of the electro-mechanics with elevation, two check stability of gap contamination. The test showed an electro-mechanical instability of the actuator 255 possibly due to the shortening of the capacitive sensor contact to the actuator body (ground). Tomorrow miniskirt will be dismounted to check and try a fixing of this actuator.


Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback